### v2

# Pulse-PPG: An Open-Source Field-Trained PPG Foundation Model for Wearable Applications Across Lab and Field Settings

Abstract.
Photoplethysmography (PPG)-based foundation models are gaining traction due to the widespread use of PPG in biosignal monitoring and their potential to track diverse health indicators. In this paper, we introduce Pulse-PPG, an open-source PPG foundation model trained exclusively on raw PPG data collected over a 100-day field study with 120 participants. Existing open-source PPG foundation models are trained on clinical data, and those trained on field data are closed source, limiting their applicability in real-world settings. Extensive evaluations demonstrate that Pulse-PPG, trained on uncurated field data, exhibits superior generalization and performance across clinical and mobile health applications in both lab and field settings, when compared with state-of-the-art PPG foundation models trained on clinical data. Exposure to real-world variability in field-collected PPG data enables Pulse-PPG to learn more robust representations. Furthermore, pre-training Pulse-PPG on field data outperforms its own pre-training on clinical data in many tasks, reinforcing the importance of training on real-world datasets. To encourage further advancements in robust PPG modeling, we have open-sourced2our Pulse-PPG model, providing researchers with a valuable resource for developing the next generation of task-specific PPG-based models.
Foundation models, Wearables, Health-wellbeing, Contrastive Learning
$\bm{\dagger}$Both co-first authors contributed equally to this research.
copyright: acmlicensedjournal: IMWUTjournalyear: 2025journalvolume: 9journalnumber: 3article: 126publicationmonth: 9doi: 10.1145/3749494ccs: Human-centered computing Ubiquitous and mobile computingccs: Computing methodologies Machine learning
1. Introduction
Unobtrusive, ubiquitous, and cost-effective wearable sensors have demonstrated the potential to revolutionize real-time monitoring of health and wellness by enabling the detection of various physical and mental health states. Photoplethysmography (PPG) in smartwatches has emerged as a widely used modality due to its non-invasive assessment of physiology without the need for firm attachment. It is used for estimating physiological metrics such as heart rate, heart rate variability (Sarhaddi et al., 2022), blood glucose (Ali et al., 2024), oxygen saturation (Rajakariar et al., 2024), and blood pressure (He et al., 2022; Fortino and Giampà, 2010). For diagnosis, it can detect cardiovascular conditions (Ouyang et al., 2017), including atrial fibrillation (Poh et al., 2018) and detect hypoxia (Lazazzera et al., 2020). For mental health and wellness, it can track stress (Zhu et al., 2023), emotion (Kontaxis et al., 2020), focus (Wang et al., 2024), and depression (Kontaxis et al., 2020).
However, PPG-based inference in real-world settings remains challenging due to its high susceptibility to noise from motion artifacts (Pollreisz and TaheriNejad, 2022), ambient light (Cubas and Prado, 2023), and skin conditions (Ajmal et al., 2021). This has slowed the progress in realizing the full potential of PPG in the natural environment. For example, high accuracies for stress classification has been reported on lab data (Choi et al., 2022; Hasanpoor et al., 2022; Motaman et al., 2022; Alshareef et al., 2022; Mitro et al., 2023), but they do not generalize to the field settings (Mitro et al., 2023). Training models on lab data from the same participants and then applying it to their field data can lead to a better performance (Toshnazarov et al., 2024), but it doesn’t scale to unseen participants. Models trained using larger field datasets report low performance (Zhang et al., 2024a; Aqajari et al., 2024).
Figure 1. Overview of our open-sourced222 Pulse-PPG model weights and code repository are available at https://github.com/maxxu05/pulseppg. Pulse-PPG Foundation Model. Our model is trained on wearable field PPG with relative contrastive learning, based on the relative distances captured in our learned motif-based distance function. This model demonstrates strong performance on a wide variety of tasks across wearable field, wearable lab, and clinical settings.
The emergence of Foundation Models (FM) offers a new opportunity to address these challenges and accelerate our progress. In other domains such as natural language processing and computer vision, the foundation model paradigm has transformed the development of machine learning solutions to real-world problems (App, 2025; Met, 2025). The key property of an FM is that it is pre-trained on a large-scale dataset to ensure that the resulting feature representation encompasses all of the complexity of the data domain, which is then validated by demonstrating that the FM can solve multiple downstream tasks without additional representation learning. In computer vision, the field has moved away from collecting individual special-purpose datasets and training task-specific models to leveraging existing foundation model representations, such as DINOv2 (Oquab et al., 2023) in solving a variety of perceptual tasks. A key enabling property is that while the training datasets are private and not publicly available, the model weights are released to the research community, enabling everyone to benefit from its powerful representation. This transition in utilizing publicly available FMs has not yet occurred for mHealth and is a crucial next step.
While there have been some exciting recent efforts to develop FMs for PPG (Abbaspourazad et al., 2023; Pillai et al., 2024), current approaches suffer from two important limitations: 1) Private models which define the SOTA but whose weights are not available to the research commmunity (Abbaspourazad et al., 2023), and 2) Models which demonstrate impressive performance on lab-collected data but have not been developed or evaluated for use in the field environment (Pillai et al., 2024). The growing availability of field datasets has created tremendous opportunities for advancing PPG-based research. However, many of these datasets (Truslow et al., 2024; van Gilst et al., 2019; Schmidt et al., 2019) remain inaccessible or the computational resources required to fully exploit their potential are prohibitively expensive for many researchers and practitioners.
In this paper, we present an open-source2 foundation model for wearable field PPG signals, Pulse-PPG, which can be seen in Figure 1. We pre-train our model on the raw PPG signals from a large-scale wearable field dataset, composed of approximately 200 million seconds of raw PPG data collected from 120 participants. We avoid filtering our signals because such real-world noise patterns offer meaningful contextual cues on a given individual’s motion or environment, which we show enables better performance on wearable field, lab, and clinical downstream tasks, compared to training on cleaner clinical data. To address the unique challenges of field-collected PPG data, we adapted a motif-based relative contrastive learning framework (Xu et al., 2024a, c) to flexibly handle the temporal and noise characteristics of the field PPG signals and extract subtle but semantically meaningful patterns. Our empirical evaluation demonstrates that this approach achieves state-of-the-art performance, outperforming a recent PPG foundation model on 10/11 of our downstream evaluation tasks across five different datasets, covering wearable data from the field, wearable data from the lab, and clinical PPG domains.
With the weights publicly available, researchers can build upon our work for their own specialized tasks and bypass the costly and time-consuming process of training from scratch, significantly lowering the barrier to entry for those without access to extensive computational resources. Moreover, our model can serve as an embedding function, enabling researchers to extract meaningful physiological representations from raw PPG data. Beyond these, the availability of our model weights will encourage reproducibility and benchmarking across datasets and settings. Researchers can compare their methods against a common baseline, accelerating research progress.
The key contributions of this work are as follows:
  1. (1)
    We present Pulse-PPG, the best-performing open-source2 PPG foundation model trained entirely on raw, uncurated PPG data collected from a 100-day field study involving 120 subjects from diverse backgrounds.
  2. (2)
    We show that training a motif-based representation learning framework with relative contrastive learning outperforms the current state-of-the-art open-source PPG foundation model (Pillai et al., 2024) across a wide set of downstream datasets and tasks for wearable field and clinical applications. We also show that our Pulse-PPG model with 28.5 million parameters outperforms substantially more powerful general-purpose time-series models—Chronos with 200 million parameters (Ansari et al., 2024) and MOMENT with 385 million parameters (Goswami et al., 2024).
  3. (3)
    We show that training on field data achieves better performance for two different foundation model approaches than a lab-trained model on wearable lab downstream tasks, highlighting the field-to-lab generalizability of the models, with additional potential implications for clinical settings.
2. Related Works
2.1. Foundation Models for PPG signals
There have been some recent work on PPG-specific foundation models, but none of them are open and designed for wearable PPG signals (i.e., collected from a smartwatch). The closest work with open model availability is PaPaGei, an open-source PPG foundation model that improves classification and regression performance across 14 tasks (Pillai et al., 2024), including some wearable lab tasks. However, it is exclusively trained on clean clinical PPG signals, and was not evaluated on field data (Pillai et al., 2024). We show (Section 6.1) that training on this clinical data negatively impacted its generalization performance on wearable tasks involving field-collected data. The most relevant work for wearable PPGs is a foundation model pre-trained on a large-scale wearable field PPG dataset and presented in (Abbaspourazad et al., 2023). However, the model is closed-source and its training datasets are private, restricting accessibility for the research community. Similarly, SiamQuality is a foundation model pre-trained on clinical data that demonstrates generalizability on wearable lab and clinical datasets, but the model remains private (Ding et al., 2024).
Additional related works addressed tasks and goals that are adjacent to this paper. REGLE employs autoencoders to extract disentangled embeddings from clinical PPGs, focusing on their applicability to genetic and disease prediction tasks (Yun et al., 2024). TS2TC introduces a generative self-supervised learning framework for clinical PPG data but does not provide pre-trained weights and has been evaluated exclusively on clinical PPG regression tasks (Zhang et al., 2024b). PPG-PT demonstrates the utility of pre-training, but does not meet all of the requirements of a foundation model, as the model was only evaluated on one clinical PPG-specific downstream task (Davies et al., 2024).
While these prior works have made significant strides in PPG representation learning, they either lack open-source accessibility, generalizability across tasks, or applicability to field-collected wearable data, highlighting a critical gap that our work aims to address. With the release of our foundation model’s weights, ours will be the first open-source2 PPG model trained exclusively on field-collected data.
2.2. Foundation Models in General
Our work is inspired by prior successes in developing foundation models for other data modalities. These works have shown that models pre-trained in large-scale datasets can capture underlying data representations and are highly adaptable for various downstream applications (Bommasani et al., 2021). These models have had a substantial impact in fields like Natural Language Processing and Computer Vision. In the language domain, models such as BERT (Devlin et al., 2018) and the Generative Pretrained Transformer (GPT) (Radford et al., 2018) have demonstrated remarkable generalizability across a diverse range of natural language processing tasks, catalyzing the development of large language models (i.e., ChatGPT (Zhou et al., 2024)). The vision domain has seen similar breakthroughs with models like CLIP (Radford et al., 2021) and the Segment Anything Model (Kirillov et al., 2023) that are trained on large-scale data and exhibit strong performance on a broad set of downstream tasks. These works have demonstrated that pre-training on large-scale data can yield effective feature representations and serve as the motivation for this work.
3. Pulse-PPG Methodology
Our goal is to develop a foundation model capable of learning a generalizable embedding space that can be effectively applied to a diverse set of datasets (ranging from clinical PPG to mHealth field PPG) and downstream tasks (ranging from blood pressure prediction to stress detection). In Section 3.1, we detail our rationale for pre-training the foundation model on noisy field PPG data, as opposed to the traditional approach of pre-training on clean, lab-study-based PPG data. Then, in Section 3.2, we outline the key challenges involved in learning a robust foundation model in this setting and describe how our methodology addresses each of these challenges.
3.1. Motivation for Pre-training on Field PPG data
Traditionally, prior work has focused on training models using clean, lab-collected PPG data and evaluating them on mHealth datasets (Pillai et al., 2024). This approach is largely due to the significant challenges associated with utilizing real-world field data, which is inherently noisy and variable. However, we hypothesize that by designing a methodology capable of effectively learning patterns from field data, we can develop more robust representations, generalize better, and achieve superior performance on downstream mHealth tasks, such as instantaneous heart rate prediction or activity recognition. In this way, the noise present from the wearable field sensor becomes a benefit, as it could introduce additional contextual information of a person’s current state and environment. Importantly, pre-training on mHealth data and evaluating on mHealth data reduces domain shift, as the model is exposed to the same types of noise, variability, and environmental factors during both training and evaluation.
Figure 2. Random 5s Snippets of PPG signals from Clinical, Wearable Lab, and Wearable Field Settings. Signal quality declines gradually but manageably when moving from clinical-grade PPG in hospital settings to wearable PPG in controlled lab conditions. However, a marked deterioration occurs in the uncontrolled real-world environment of the wearable field setting, driven by daily wear factors, such as motion artifacts (Pollreisz and TaheriNejad, 2022), ambient light (Cubas and Prado, 2023), and poor sensor contact (Lo Grasso et al., 2024).
Clinical datasets, while less noisy and more consistent due to their controlled collection processes administered by professionals, lack the dynamic variability and environmental factors present in real-world settings. For example, clinical datasets often involve participants with specific health conditions and are collected under static, controlled conditions, which fail to capture the wide range of physiological and environmental variability encountered in the wild. In contrast, field data includes participants with and without specific conditions, as well as a diverse array of physiological patterns influenced by real-world factors, such as poor sensor contact from loose bands and daily motion. This richness and diversity make field data a more comprehensive source for pre-training a foundation model.
By pre-training on field data, we aim to create a foundation model that better reflects the complexity of real-world scenarios and generalizes more effectively across diverse applications.
3.2. Learning Robust PPG Representations from Field Data
To build a robust foundation model, we must design a pre-training task that not only captures high-level, obvious semantic features, such as heart rate changes, but also uncovers subtle, latent patterns in the data, such as variations in pulse waveform morphology or transient physiological responses. PPG signals, which measure blood volume changes, encode a wealth of physiological information. However, this information is often obscured by noise, particularly in the field. Traditional approaches address noise by normalizing or aggregating data into statistical features and training models on these features. While this reduces the impact of noise, it often discards fine-grained details that are critical for accurate inference or discovering hidden clusters and trends in the data.
To overcome these challenges, we introduce Pulse-PPG, a method specifically designed to address three key issues: 1) the development of a domain-specific pre-training task that leverages PPG-specific domain knowledge; 2) exploiting the inherent noise of real-world PPG data; and 3) the ability to capture subtle, yet meaningful, patterns in the data that are essential for ensuring generalizability across diverse and dynamic real-world settings.
3.2.1. PPG-Specific Pre-training Task
A classic self-supervised objective for time-series foundation models is contrastive learning, where augmentations (e.g., cropping, scaling, jittering, shuffling) create distinct but semantically similar positive pairs (Meng et al., 2023; Yue et al., 2022; Zhang et al., 2022; Eldele et al., 2023; Woo et al., 2022; Yang and Hong, 2022; Yang et al., 2022; Ozyurt et al., 2022; Lee et al., 2022). However, PPG signals face unique challenges: (1) time-series lack image-like invariances (e.g., rotation), and (2) augmentations can distort PPG semantics. For example, scaling alters amplitude (critical for blood volume inference) and shuffling disrupts temporal dependencies (key for heart rate variability). Alternatively, subject-aware contrastive learning avoids augmentations by leveraging long-term recordings from individual subjects (Abbaspourazad et al., 2023). While especially effective for biosignals (Jeong et al., 2024), existing methods rely on naive temporal sampling (Tonekaboni et al., 2021) or supplemental context (e.g. event labels) (Jeong et al., 2023). To address this, we propose directly exploiting raw PPG semantics in our large-scale unlabeled dataset, eliminating reliance on augmentations or auxiliary data. To unlock the potential of our large-scale, unlabeled wearable field PPG dataset, we need a methodology that utilizes semantic information from only the raw PPG signals.
A critical characteristic of raw PPG signals lies in their origins in the cardiopulmonary system. PPG signals are pulsative: quasiperiodic and composed of repeating motifs that correspond to individual cardiac cycles (Allen, 2007; Elgendi, 2012). We define motifs abstractly as the short temporal shapes within the overall PPG morphology that reflect specific semantic information, such as the first upward slope in a specific PPG beat that represents systolic rise time. These motifs encode critical physiological information, such as heart rate variability, blood pressure changes, and vascular dynamics.
Therefore, we propose to utilize our pre-training task to learn differences between PPG instances, based upon the differences in their motifs, drawing from Xu et al. (2024b). In this way, the distance function, $d(\textbf{X}_{\text{anchor}},\textbf{X}_{\text{cand}})$ can compare two PPG sequences, $\textbf{X}_{\text{anchor}}\in\mathbb{R}^{T}$ and $\textbf{X}_{\text{cand}}\in\mathbb{R}^{T}$ by finding the motif in $\textbf{X}_{\text{cand}}$ that best matches a given motif $\textbf{X}_{\text{anchor}}$. This idea can be seen below in Equation 1, with the ”matching” function given by a distance function $d_{m}$.
(1) $d(\textbf{X}_{\text{anchor}},\textbf{X}_{\text{cand}})\approx\sum_{i=0}^{T}\underset{j\in[0,\cdots,T]}{\text{argmin }}\Big{\{}d_{m}\Big{(}\textrm{Motif}\big{(}\textbf{X}_{\text{anchor}}[i]\big{)},\textrm{Motif}\big{(}\textbf{X}_{\text{cand}}[j]\big{)}\Big{)}\Big{\}}$
This distance function would enable semantically aware comparisons between PPG sequences, learning an embedding space that captures physiologically meaningful relationships.
3.2.2. Exploiting the Noise of Real-World, mHealth PPG data
Many prior works that utilize morphology information from raw PPG signals for their task-specific application, such as for blood pressure prediction (Salah et al., 2022; Shin and Min, 2017) or stress detection (Ahmed et al., 2022; Rinkevičius et al., 2019), will first utilize a peak detection algorithm to segment individual PPG beats that correspond to as single cardiac cycle (Haddad et al., 2020; Kavsaoğlu et al., 2016; Fischer et al., 2016). After segmentation, specific temporal features within the beat can be investigated, such as the dicrotic notch or the systolic rise time. However, these methods are evaluated exclusively on clinical or lab-collected PPG data, where signal quality is controlled and high.
Exploiting PPG morphological information is significantly more challenging in real-world wearable field PPG data, where persistent noise and motion artifacts make beat-to-beat segmentation and morphological feature extraction infeasible. As shown in Figure 2, the sharp decline in PPG signal quality, ubiquitous in mobile health settings, severely undermines the reliability of traditional segmentation and feature isolation methods.
To address this challenge, we utilize a learnable motif-based distance function (Equation 2) trained directly on noisy wearable field PPG data. Unlike traditional approaches that depend on predefined morphological features or clean beat segmentation, our method learns to identify motifs from raw, uncurated signals. By training under real-world noise conditions, we leverage the inherent diversity of field data motifs. Crucially, noise patterns are not discarded but instead leveraged as meaningful contextual cues: non-random artifacts (e.g., from motion or environment) correlate with specific activities, states, or conditions, enhancing motif uniqueness. This enables the model to identify semantic clusters and learn robust representations generalizable to diverse real-world scenarios.
(2) $d(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{cand}})$ $\coloneqq\sum_{\textbf{x}_{\textrm{anc}}\in_{s}{\textbf{X}_{\textrm{anc}}}}\Bigg{(}\bigg{(}\sum_{\textbf{x}_{\textrm{cand}}\in_{s}{\textbf{X}_{\textrm{cand}}}}\underset{{\textbf{x}_{\textrm{cand}}\in_{s}{\textbf{X}_{\textrm{cand}}}}}{\textrm{softmax}}\Big{(}\textrm{sim}\big{(}f_{q}({\textbf{x}}_{\textrm{anc}}),f_{k}(\textbf{x}_{\textrm{cand}})\big{)}\Big{)}f_{v}(\textbf{x}_{\textrm{cand}})\bigg{)}-\textbf{x}_{\textrm{anc}}\Bigg{)}^{2}$
(3) $f_{\{q/k/v\}}({\textbf{X}})$ $=\textrm{DilatedConvNet}_{\{q/k/v\}}\big{(}{{\textbf{X}}}\big{)}$
where $\smash{\textbf{X}\in\mathbb{R}^{T\times D}}$ and $\smash{\textbf{x}\in\mathbb{R}^{D}}$ with $T$ as the time length, $\in_{S}$ as a $\in$ but with a subsampling of stride $s$, and $sim()$ is a dot product. This function is implemented with a standard cross-attention block (Vaswani et al., 2017) with projection and normalization layers being omitted for brevity.
This trainable function closely matches our idea presented in Equation 1. $d_{m}$ is modeled as a reconstruction error, where the motifs of $\textbf{X}_{\text{cand}}$ that best matches a given motif $\textbf{x}_{\text{anchor}}$ are used to reconstruct $\textbf{x}_{\text{anchor}}$, in a kernel regression. Motifs are captured with the dilated convolutions of $f_{\{q/k/v\}}$, as seen in Equation 3. The weights in the kernel regression found in $\smash{(\sum_{{\textbf{X}_{\textrm{cand}}}}{\textrm{softmax}})}$ serve the purpose of the original argmin, where the function is learned to find the weights that minimize the reconstruction error.
This work builds on Xu et al. (2024a), which proposed a proof-of-concept motif-based distance for physiological signals in controlled, small-scale settings. However, their approach was limited to pre-training and evaluation on the same small labeled dataset, lacking the scalability and generalizability required for real-world foundation models. For example, the original work only evaluated on sequences with a maximum length of 3,840 time-points, much shorter than our 4-minute-long PPG input lengths of 12,000. This is significant because the backbone of the training function utilizes a cross-attention function, with a computational complexity of $O(T^{2})$, but we were able to address this with our $\in_{s}$ function with $s=10$. This allows for the reconstruction error to be calculated from subsample of the original sequence, effectively reducing computational complexity to $O((T/s)^{2})$, but without degrading the quality of motif similarity comparisons by allowing $f_{\{q/k/v\}}$ to capture the motifs found from the full sequence. Additionally, the prior work (Xu et al., 2024a) utilized a simple, untuned dilated convolution encoder to act as a control to fairly compare self-supervised learning methods. In our work, we utilize a much more complex ResNet encoder backbone to develop a powerful foundation model that can capture and handle the noise and diversity in real-world PPG data. To this end, it necessitated an extensive tuning of the methodology (see Section 4.3.1) to work in our setting, but due to this, we have been able to demonstrate robustness in the dynamic field environment across diverse downstream tasks and datasets.
3.2.3. Capturing Subtle Yet Meaningful Patterns
Given our motif-based distance function, we can identify positive and negative pairs to draw together and push apart in the embedding space, respectively. Traditionally, contrastive approaches will utilize the Normalized-Temperature Cross-Entropy Loss (NT-Xent) function (Equation 4), in which the positive pair is pulled towards the anchor, and all negatives are pushed away equally.
(4) $\ell(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{pos}},\mathcal{S}_{\textrm{neg}})=-\log\frac{\exp(\textrm{sim}(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{pos}})/\tau)}{\sum_{\textbf{X}_{\textrm{neg}}\in\mathcal{S}_{\textrm{neg}}}\exp(\textrm{sim}(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{neg}})/\tau)+\exp(\textrm{sim}(\textbf{X}_{\textrm{anc}},\textbf{X}_{\textrm{pos}})/\tau)}$
The standard NT-Xent approach’s reliance on rigid positive/negative assignments makes it prone to false positives (misclassifying dissimilar instances as positives) and false negatives (overlooking valid positives in the candidate pool). For example, in stress assessment, being stressed due to bad traffic differs from the experience of being tailgated, but the two are more similar compared to being engaged in a verbal conflict. Treating all negatives equally ignores such nuances, resulting in overly coarse clusters that fail to capture subtle semantic relationships.
An alternative idea to address this is utilizing a metric learning loss function (Kim et al., 2019; Hoffmann et al., 2022; Kan et al., 2021), in which the encoder learns to embed the PPG instances exactly as described by our motif-based distance function. However, although our learned motif-based distance function will be able to capture semantic information, it was trained without labels and is not expected to always perfectly identify the correct semantic relationships.
A ”goldilocks” approach between these two ideas would be using a relative contrastive learning framework that models the relative positions of the distances. As such, we utilize the concept of relative dissimilarity via multiple negative sets (Xu et al., 2024c). For each positive pairing, a negative set is created by selecting samples whose distance from the anchor exceeds the distance of the positive pair. Each candidate in the dataset is iteratively used to form a positive pair, while the remaining candidates contribute to the negative set:
$f_{\textrm{neg}}({X}_{\textrm{anc}},{X}_{\textrm{pos}},\mathcal{S})=\{{X}\in\mathcal{S}\mid d({X}_{\textrm{anc}},{X})>d({X}_{\textrm{anc}},{X}_{\textrm{pos}})\}$
This concept is incorporated into the Relative Contrastive Loss function, which builds representation spaces that are aware of relative distances and semantic nuances:
(5) $\mathcal{L}_{\textrm{RelCon}}=\sum_{{X}_{\textrm{i}}\in\mathcal{S}_{\textrm{cand}}}\ell({X}_{\textrm{anc}},\ {X}_{\textrm{pos}}\coloneqq{X}_{\textrm{i}},\ \mathcal{S}_{\textrm{neg}}\coloneqq f_{\textrm{neg}}({X}_{\textrm{anc}},{X}_{\textrm{i}},\mathcal{S}_{\textrm{cand}}))$
By integrating relative dissimilarity into contrastive learning, the model forms tight clusters in a representation space reflecting fine-grained relationships. This improves generalization to unseen users by capturing within- and between-class nuances, ensuring robustness to intra-user variability and adaptability to new users’ physiological patterns. The resulting representations reliably handle field data variability for users unseen during training.
This work builds on Xu et al. (2024c), which proposed relative contrastive learning to learn a foundation model for accelerometry sensor data. However, accelerometry sensor data varies significantly from PPG sensor data. For example, accelerometry has a well-established set of invariances (i.e., 3D rotation) (Tang et al., 2020), due to how inertial its measurements are fundamentally governed by physics—the magnitude and temporal patterns of acceleration reflect biomechanical forces generated by movement, regardless of sensor orientation or placement. PPG signals, by contrast, exhibit no such geometric invariances—their morphology is highly sensitive to sensor contact, motion artifacts, and physiological variability. Rather than presupposing geometric symmetries, for Pulse-PPG, we allow the motif-based distance function to implicitly learn noise-robust representations directly from the raw PPG waveform’s temporal structure. Additionally, their work, as well as other accelerometry machine learning work, demonstrates that short time windows can capture semantically meaningful motion information, and as such, they only encode accelerometry segments spanning 2.56 seconds, or 256 time points. However, for Pulse-PPG, we seek to model 4-minute (12,000 time point) PPG segments, to align with prior stress models (Lamichhane et al., 2017; Wilson et al., 2018; Bari et al., 2020). Therefore, this also results in an extensive tuning of the methodology (see Section 4.3.1) to work in our setting. For example, we reduce the same-subject candidates from 20 to 1 and constrain them to be drawn from the same hour as the anchor. Due to this tuning, we have been able to develop Pulse-PPG, a powerful foundation model that generalizes across many different PPG settings and tasks.
4. Pre-training Procedure for Pulse-PPG
In this section, we provide the technical details of our foundation model’s implementation and pre-training. Section 4.1 outlines how we employ our learnable motif-based distance function and relative contrastive loss to train the Pulse-PPG foundation model. Section 4.2 describes the composition and characteristics of our pre-training dataset, detailing how we process the wearable field PPG data to align with our pre-training task, as well as highlighting the dataset’s inherent properties that make it well-suited for learning robust representations in noisy, real-world conditions. Finally, Section 4.3 describes specific implementation details, such as specific architecture and total GPU compute used. Please refer to Appendix for specific hyperparameter details.
4.1. Model Pre-training Details
4.1.1. Training the Learnable Motif-based Distance Function
As mentioned previously in Section 3.2.2, making our distance function learnable is a key strength of our approach, as it enables motif-comparisons between PPG sequences, even when such PPGs are afflicted with noise. Therefore, before utilizing our relative contrastive loss, we learn our distance function in an unsupervised fashion.
In our initial label-free pre-training setting, we do not know if a random pair of PPG sequences shares class labels, and thus, we do not know if reconstruction error should be minimized to learn a motif-matching similarity function between them. Therefore, during training, we set ${\textbf{X}}_{q}$ and ${\textbf{X}}_{k}$ to be the same value and apply a missingness mask on ${\textbf{X}}_{q}$. This mask has a contiguous segment of missingness that stretches 2 seconds. In this way, our distance function will learn to retrieve the regions from the key $\textbf{X}_{k}$ that match the missing 2-second motif from the query, $\bar{\textbf{X}}_{q}$, for reconstruction. Then, after training, we can utilize this distance as a static function for the RelCon loss to identify relative distances between ${\textbf{X}}_{q}$ and ${\textbf{X}}_{k}$, where ${\textbf{X}}_{q}$ and ${\textbf{X}}_{k}$ are different instances with different values.
4.1.2. Training the Pulse-PPG Foundation Model
Now, once we have our trained motif-based distance function, we can use it to identify relative distances of a candidate set of PPG sequences from an anchor PPG sequence. This candidate set of PPG sequences is sampled from two sources: within-subject and between-subject.
  • Sampling within-subject, across-time, draws candidate PPG sequences from the same subject, but at a different time window from the anchor. This allows us to model how an individual’s behavioral patterns throughout their day affect their PPG signal, for example, via exercise.
  • Sampling between-subject draws candidate PPG sequences from other persons within a batch. This allows us to model similarities and differences in physiology among different individuals.
Given this candidate set, we can sort the candidates with our motif-based distance function found in Equation 2 and then apply our RelCon loss function in Equation 5 based on their relative distances from the anchor. As such, our Pulse-PPG foundation model will be able to capture both within-subject and between-subject semantic information within the embedding space.
4.2. Pre-training on a Field Wearable PPG Dataset
4.2.1. Dataset Description:
For our pre-training, we obtained access to the Mobile Open Observation of Daily Stressors (MOODS) study dataset (Neupane et al., 2024) that is composed of 822,247 unique 4-minute 50 Hz PPG segments, with 122 participants who wore a smartwatch every day for up to 100 days. Participants included 39 men, 77 women, and 6 non-binary individuals, with an average age of 38 years ($SD=13$). Seventy-one were working professionals across fields such as academia, information technology, administration, public safety, health, law, sales, and construction; fifty were students pursuing higher education, and one participant did not disclose their occupation. Participants wore a study-provided Fossil Sport (Version 4) smartwatch, and approximately 55K hours of raw PPG data, sampled at 100 HZ, were collected during the study. The study was conducted using a WearOS app for smartwatches, a cross-platform smartphone app, and cloud services. Detailed descriptions of the system and study design of the MOODS study are available in (Neupane et al., 2024).
4.2.2. Subject-wise Train/Val/Test Splits:
To pre-train a model capable of handling substantial within- and between-subject variability, we randomly shuffled participants and created a train/val/test split of 84/18/18. During pre-training, the test set is never used, however, we establish this split so that the downstream tasks that utilize this same dataset (i.e., Wearable Field Stress and Activity) have a held-out test set never exposed during pre-training. We deliberately adopted a subject-based splitting strategy to pose a more challenging task for the model: making inferences on data from entirely unseen individuals, a common scenario in real-world applications.
4.2.3. Preserving Noisiness:
To pre-train the physiological model, we utilize raw photoplethysmography (PPG) data collected from the participants. Out of 122 participants, the dataset recorded PPG data from 120 participants, which we use for pre-training. PPG data collected in the field is noisy; however, avoiding signal-specific filtering can provide certain advantages for generalized, robust representations, e.g., adapting to dataset diversity, preserving subtle meaningful variations, and better transferability. Hence, we refrain from applying any particular filtering technique, but focus on addressing between-person variability through a global person-specific $z$-normalization. Please see Appendix A.3 for our ablation study, which shows that normalization consistently enhances performance.
4.2.4. PPG Sequence Inputs for Pre-Training Model
In our modeling approach, we use 4-minute long windows as sequence inputs for the Pulse-PPG Foundation Model. This choice is grounded in research highlighting the cyclical nature of physiological responses to stress. Studies have shown that stress-induced patterns in heart rate and heart rate variability (HRV) follow cyclical behaviors, with durations typically ranging from 3.5 to 4.2 minutes (Lamichhane et al., 2017; Wilson et al., 2018; Bari et al., 2020). For example, heart rate increases during stress and remains elevated until the stressor ends, after which it returns to baseline, forming a cyclical pattern. Field studies involving stressful conversations or work-related stress, have further validated these patterns, with median cycle durations around 4 minutes (Bari et al., 2020).
By segmenting PPG signals into non-overlapping 4-minute windows, we aim to capture both micro-level perturbations (e.g., transient changes in pulse waveform morphology) and macro-level cyclical patterns (e.g., sustained heart rate elevation during stress). This approach allows our model to learn representations that are sensitive to the temporal dynamics of stress, which are critical for accurate stress detection. The 4-minute window size strikes a balance between capturing meaningful cyclical patterns and retaining fine-grained temporal details. Please refer to Appendix A.3 for the ablation study of evaluating the effects of varying input window sizes. Overall, the 4-minute window yields the best performance, consistently outperforming the 1- and 2-minute models across most metrics and ranking within the top two.
It is important to note that, although our initial pre-training inputs are 4 minutes long, our model incorporates a temporal pooling mechanism that enables it to generalize to inputs of varying lengths. This flexibility allows our model to be utilized on our diverse downstream tasks, in which the input windows are variable in length.
4.3. Pulse-PPG Model Implementation Details
4.3.1. Model Adaptation and Tuning
As described in Sections 3.2.2 and 3.2.3, we adapt the recent motif-similarity-based reconstruction and relative contrastive loss-based self-supervised modeling approach (Xu et al., 2024b, c) so that it can capture the intricacies of 4-minute segments of PPG data.
We began with a parameter search for the distance function model, which was responsible for motif-based reconstruction of masked PPG segments. In this phase, we experimented with different kernel sizes (5, 11, 15, 21, 60, 120), embedding dimensions (32, 64, 128), double receptive field values (to control how many layers and how much temporal context the network could attend to), and learning rates (0.001, 0.0001). Smaller kernel sizes and shallower receptive fields failed to capture sufficient temporal context for motif alignment, while overly large fields led to over-smoothing and poor reconstruction on noisy segments. The selected configuration provided the best trade-off between learning localized motifs and capturing broader temporal dependencies in field data.
Once the distance model was trained, we used it to train a 1D ResNet self-supervised encoder. In our parameter exploration for this ResNet encoder, we evaluated different kernel sizes (9, 11, 15), initial base filter sizes (32, 64, 128), stride sizes (1, 2, 5, 10, 15, 20), number of ResNet blocks (4, 8, 12, 16), number of within-user same-hour candidates (1, 5, 10, 15), final pooling strategies (mean, max), and learning rates (0.001, 0.0001). We found that while deeper architectures marginally improved performance on training subsets, they generalized poorly on held-out field data, likely due to overfitting on participant-specific noise. Pooling strategies also played a significant role: max pooling consistently produced more discriminative embeddings for downstream tasks, especially when separating subtle variations in physiological activity. Once we identified the best combination that gave us reasonable performance on the 10-day dataset, we initiated the same process for the full 100-day dataset, using a more informed combination of parameters rather than a random search. In this stage, we repeated the parameter search process for both learning stages and found the best hyperparameter combinations to be:
  • Distance learning: kernel size = 15, embedding dim = 64, double receptive field = 5, epochs = 20, learning rate = 0.001, batch size = 16, and stride = 10.
  • Contrastive learning: base filters = 128, kernel size = 11, stride = 2, groups = 1, nblocks = 12, final pool = max, epochs = 20, learning rate = 0.0001, and batch size = 64.
4.3.2. Encoder Details:
We utilize a 1D ResNet-26 network from the https://github.com/hsd1503/resnet1d code repository as our encoder backbone, with a total of 28.5m trainable parameters. There is instance normalization placed at the start of the ResNet to account for PPG distribution shifts across users and datasets. Our ResNet implementation utilizes an initial filter size of 128, a kernel size of 11, and a stride of 2, with 12 residual blocks that linearly increase the filter size every 4 blocks. There is a global average temporal pooling at the end of the network for our model to generalize across different time scales. As a result of this, our encoder embeds variable-length PPG signals into a single 1D 512-dimensional embedding vector.
4.3.3. Distance Function Details:
We utilize the code from the https://github.com/maxxu05/rebar code repository to implement the distance function. This is a lightweight model, used as a static function to identify relative distances for the RelCon loss function, and only has 127k trainable parameters. The dilated convolution network is a series of dilated convolution blocks, made up of a dilated convolution layer with a residual connection skip, followed by a ReLU and instance norm. There are 5 blocks, with an initial dilation of 1 that doubles after each layer. The filter size is set to 64, the kernel size is set to 15, and the stride is set to 10. The input convolution layer to the DilatedConvNet is set to be a partial convolution (Liu et al., 2018) to handle missingness in our pre-training and application procedures. The reconstruction error used as our distance function is a standard MSE loss.
4.3.4. Training Details:
Our Pulse-PPG foundation model was trained for 5 days for 6 epochs on an NVIDIA L40S GPU. Each epoch is composed of 606,833 unique 4-minute PPG segments or about 5,424 unique participant days, and due to this large data size, each epoch took 20 hours to finish. We chose this early stopping point due to convergence of the pre-training loss, (2) diminishing returns on downstream task performance, and (3) computational efficiency—extending further would yield negligible benefits while significantly increasing the carbon footprint. This can be seen in the training and validation loss curves visualized in Figure 5. Both training and validation loss are calculated for each batch, but aggregated across the entire epoch for reporting. The training loss is backpropagated and weights are updated after each step for each batch (accelerating convergence and lowering computational cost), aligning with deep learning standards (Amari, 1993). An Adam optimizer was used with a learning rate set to 0.0001, betas set to (0.9, 0.999), no weight decay, and a batch size of 64.
The distance function was trained over 10 hours for 20 epochs with an Adam optimizer with a learning rate of 0.001, betas set to (0.9, 0.999), no weight decay, and a batch size of 16. Models were constructed with the PyTorch Python package, and full reproducibility details can be found in our final released codebase.
5. Evaluation Set-up
The power of a foundation model lies in its ability to generalize across a broad range of tasks and datasets. This makes them especially valuable to the research community, providing pre-trained, general-purpose representations that can be easily adapted to new tasks, streamlining research efforts without the need for extensive task-specific data collection or manual feature engineering. Therefore, we design our experiments to validate this by evaluating our Pulse-PPG Foundation Model on 11 different downstream tasks, across 5 datasets, spanning typical wearable PPG tasks (i.e., stress classification) and typical hospital PPG tasks (i.e., systolic blood pressure regression). In the following sections, we will describe our evaluation set-up and experimental results. We describe our downstream datasets and tasks in Section 5.1, evaluation methodologies in Section 5.2, evaluation metrics in Section 5.3 and experimental results in Section 6.
5.1. Downstream Datasets and Tasks
In order to evaluate the generalizability of our Pulse-PPG foundation model, we evaluate our 5 downstream datasets and 11 tasks, which can be seen in Table 1. Our datasets have a range of different ”PPG type” settings:
  • Wearable Field PPGs reflect the real-world setting for wearable applications with low signal quality.
  • Wearable Lab PPGs originate from controlled lab studies, in which subjects are given a wearable PPG sensor and asked to do a set of scripted activities under supervision. Generally, signal quality is higher.
  • Clinical PPGs originate from a hospital or another heavily monitored environment, in which PPG signals are collected from a finger sensor on stationary patients. Generally, signal quality is high.
MOODS (Neupane et al., 2024) PPG-DaLiA (Reiss et al., 2019a) WESAD (Schmidt et al., 2018a) SDB (Garde et al., 2014) PPG-BP (Elgendi et al., 2022)
PPG Type Wearable Wearable Wearable Clinical Clinical
Quality Type Field Clean Clean Clean Clean
Num of Subjects 120 15 15 146 219
Time per Subject 100 Days 103 Minutes 60 Minutes 360 Minutes 10 Seconds
Tasks Available Stress (2) Activity (2) Instant HR (R) Activity (9) Stress (4) Stress (2) Sleep Disturbance (2) Systolic BP (R) Diastolic BP (R) Avg. HR (R) Hypertension (2)
Table 1. Downstream Datasets and Tasks. $(\bm{\cdot})$ designates either a regression task or # of classes for a classification task.
Additionally, with each of the PPG types, there is a set of canonical downstream tasks associated with them. For wearables, that includes stress classification (Schmidt et al., 2018a; Neupane et al., 2024), activity classification (Neupane et al., 2024; Reiss et al., 2019a), and instantaneous HR regression (Reiss et al., 2019a). For clinical settings, this includes sleep disturbance classification (Garde et al., 2014), blood pressure regression (Elgendi et al., 2022), and hypertension  (Elgendi et al., 2022). While activity classification is traditionally performed using IMU sensors or motion signals, we include it as a downstream task to showcase the versatility of our model in learning meaningful representations solely from PPG data. Each of these canonical tasks is captured in our downstream datasets. Please see Appendix A.1 for further dataset details.
5.2. Evaluation Implementation Details
All PPG signals were resampled to match the 50 Hz sampling frequency present in our pre-training MOODS dataset (Neupane et al., 2024). For our evaluations using PaPaGei, we made sure to resample the signals to 125 Hz to match the sampling frequency that they trained on (Pillai et al., 2024).
To train and evaluate our linear probes, we combined the train and validation splits and used a cross-validation hyperparameter grid search with the Scikit Learn package. For the regression tasks, we utilized a linear regression with an L2 regularization and conducted a grid search over the following parameters: ’alpha’:[0.1,1.0,10.0,100.0],’solver’:[’auto’,’cholesky’,’sparse_cg’] with the scoring function as negative MSE. For the classification tasks, we utilized a logistic regression with an L2 regularization and lbfgs solver and conducted a grid search over the following parameters ’C’:[0.01,0.1,1,10,100],’solver’:[’lbfgs’],’max_iter’:[1000,10_000] with the scoring function as macro F1. Inputs into the probes were normalized with the Scikit Learn StandardScaler module.
In order to train and evaluate our fine-tuning results, we fit a classification or regression head on top of the existing network, then train the entire network end-to-end. For classification, the head was made up of a layer normalization layer, followed by a linear layer, with the loss function being a linear combination of Cross Entropy loss, a soft F1 loss (Bénédict et al., 2021), and Dice loss (Li et al., 2019). For regression, the head was an MLP layer, with an initial layer with an output dimensionality of 128, followed by the GeLU activation, followed by a final linear layer. The loss function used was the MSE loss function. For both evaluations, we utilize an Adam optimizer. The checkpoint with the highest macro F1 score and the lowest MAE on the validation split was selected for inference on the test set. For exact implementation details, please refer to our public codebase: https://github.com/maxxu05/pulseppg.
5.3. Evaluation Metrics
For evaluating our models across diverse downstream tasks, we utilize task-specific metrics for both regression and classification problems. For regression tasks, we report standard Mean Absolute Error (MAE) and Mean Squared Error (MSE). Additionally, we use Mean Absolute Percentage Error (MAPE) to assess relative errors across different scales. For classification tasks, we report macro F1 score, Accuracy, Precision, and Recall to evaluate overall model performance and the trade-offs between false positives and false negatives. To provide a more comprehensive assessment, particularly for imbalanced datasets, we compute Area Under the Precision-Recall Curve (AUPRC) and Area Under the Receiver Operating Characteristic Curve (AUROC), which measure the model’s ability to distinguish between classes.
6. Experimental Results
Our primary claims are that (1) Pulse-PPG is able to achieve state-of-the-art performance across a wide set of downstream PPG tasks for wearable and clinical applications, (2) pre-training on wearable field data achieves surprisingly strong performance compared to cleaner clinical data, and that (3) researchers will be able to utilize our open-sourced model in a variety of use-cases to expedite their own research efforts.
We evaluate these claims through a series of experiments: Pulse-PPG vs. PaPaGei, a prior PPG FM (Section 6.1); Pulse-PPG vs. General Time-series FMs and Supervised Baselines (Section 6.2); Assessing Field-to-Lab Generalizability (Section 6.3); and finally the Evaluation of Pulse-PPG use-cases (Section 6.4).
6.1. Experiment 1: Pulse-PPG vs. PaPaGei, A Prior PPG foundation model
6.1.1. Background
To evaluate the effectiveness of our Pulse-PPG foundation model, we benchmark against the current state-of-the-art in open PPG foundation model, PaPaGei (Pillai et al., 2024). Their PaPaGei-S model uses a 1D ResNet backbone and leverages a novel self-supervised objective with PPG-morphology specific metrics, like the Inflection Point Area ratio, achieving top performance on clinically-focused downstream tasks. Crucially, they release model weights and code (https://zenodo.org/records/13983110), which we utilize for replication.
However, their pre-training dataset was composed solely of pre-processed, clean clinical PPG signals. As previously discussed in Section 3.2.2, we hypothesize that their PPG-morphology specific metrics will be unable to generalize well in the noisier field setting. This is unlike our learned distance metric, which can potentially learn to do motif-similarity comparisons, even in noisy settings.
We also note that while PaPaGei demonstrated strong performance across various datasets in their work, their evaluation methodology involved using multiple different hyperparameter-tweaked versions of the foundation model and choosing the best model per task (Pillai et al., 2024). We report results on one set of pre-trained weights per model (for both PaPaGei and Pulse-PPG) throughout our entire paper, which explains any discrepancies in our reported numbers versus their original work. While PaPaGei sets a critical stepping stone for PPG foundation modeling, our goal is to build a single open-source2 PPG foundation model that can generalize across many tasks.
[…]
This figure compares the Pulse-PPG model with the PaPaGei foundation model, showcasing performance in classification (F1 score) and regression (MAPE). Pulse-PPG consistently outperforms PaPaGei, especially in Wearable Field and Lab PPG tasks.
Figure 3. Comparison of Pulse-PPG vs. PaPaGei (Pillai et al., 2024), a prior PPG foundation model. The two plots show the relative performance of a linear probe evaluation for each model, with the left for classification via F1 score and the right for regression via Mean Average Percentage Error. 10/11 of the task data points reside above the slope, demonstrating how Pulse-PPG consistently outperforms PaPaGei, with particularly substantial improvements for the Wearable Field and Lab PPG tasks. See further discussion in Section 6.1.3.
6.1.2. Experimental Design
To ensure a fair comparison, we utilize a linear probing evaluation for each downstream task for both Pulse-PPG and PaPaGei-S models. Linear probing is a widely accepted evaluation method for foundation models because it isolates the quality of each foundation model’s learned representations by freezing the model’s embeddings and training only a simple logistic classifier or linear regression. Both PaPaGei and Pulse-PPG embed PPG sequences into a single 1D 512-dimensional vector, making their linear probe results directly comparable. By using linear probing—training a simple classifier on frozen embeddings, we isolate the quality of the learned representations, avoiding biases from architectural differences.
To re-emphasize, this linear probing evaluation is not intended to achieve maximum task-specific performance, instead serving as a controlled experiment to enable a fair comparison between the foundation models.
6.1.3. Results
Figure 3 illustrates the difference in performance between our Pulse-PPG foundation model and the publicly released pre-trained PaPaGei. Full results can be found in Table 7 in the Appendix. Pulse-PPG outperforms PaPaGei on 10/11 of our downstream tasks, consistently across many of the reported metrics.
On the Wearable PPG tasks, our Pulse-PPG achieves a consistently large .$\geq.{\color[rgb]{0,0,0}10}$ increase in F1 scores in the classification tasks, except one. Then, on the instantaneous HR regression task, our Pulse-PPG model has a large 6.28 improvement in MAE. These heart rate estimates are conducted on unfiltered wearable PPG data, which is significant due to the widely recognized challenge of noise-induced degradation in heart rate estimation. On the Clinical PPG tasks, our Pulse-PPG model continues to outperform PaPaGei, with $\sim 0.5$ improvement in MAE for Systolic BP, Diastolic BP, and average HR regression, as well as a .05 improvement in F1 score on hypertension classification. While clinical evaluation tasks show smaller improvements compared to wearable ones, the consistent gains highlight the model’s cross-domain generalizability.
The only task on which Pulse-PPG underperforms is the Sleep Disturbance task. This can be attributed to the nature of Pulse-PPG’s pre-training dataset, MOODS, where participants were instructed to wear the smartwatch only during waking hours (Neupane et al., 2024). Consequently, Pulse-PPG was not exposed to PPG signals during sleep in its pre-training phase, and human physiology during sleep is significantly different from that of waking hours due to natural biological processes (Smets et al., 2018). In contrast, PaPaGei’s pre-training included a sleep study dataset (Chen et al., 2015).
6.2. Experiment 2: Pulse-PPG vs. General Time-series FMs and Supervised Baselines
Table 2. Pulse-PPG vs. PPG FM, General Time-series FM, and Supervised Baselines. Mean Performance across Classification and Regression Tasks. The best is bolded, and the second is underlined. Full results can be found in Table  6. Pulse-PPG achieves stronger performance across all metrics than the prior PPG foundation model, PaPaGei. It always maintains Top-2 performance, even among general time-series foundation models, which have significantly higher parameter count (Pulse-PPG with 28.5m vs. Chronos with 200m and MOMENT with 385m) and among fully-supervised models.
Avg. Classification Performance Avg. Regression Performance
Model F1 Score Accuracy Precision Recall AUPRC AUROC MAE MSE MAPE
PPG FM Pulse-PPG $\bm{{\color[rgb]{0,0,0}0.5870}\pm 0.17}$ $\bm{{\color[rgb]{0,0,0}0.6840}\pm 0.17}$ ${\color[rgb]{0,0,0}0.6237}\pm 0.16$ $\bm{{\color[rgb]{0,0,0}0.5894}\pm 0.17}$ ${\color[rgb]{0,0,0}0.5149}\pm 0.20$ $\bm{{\color[rgb]{0,0,0}0.7481}\pm 0.17}$ ${\color[rgb]{0,0,0}8.8593}\pm 6.25$ ${\color[rgb]{0,0,0}145.255}\pm 171.08$ ${\color[rgb]{0,0,0}0.1009}\pm 0.05$
PaPaGei(Pillai et al., 2024) $0.4994\pm 0.14$ $0.6361\pm 0.15$ $0.5530\pm 0.14$ $0.5210\pm 0.14$ $0.4164\pm 0.12$ $0.6840\pm 0.10$ $10.7488\pm 7.75$ $204.6825\pm 237.78$ $0.1288\pm 0.10$
General FM Chronos(Ansari et al., 2024) ${\color[rgb]{0,0,0}0.5845}\pm 0.12$ ${\color[rgb]{0,0,0}0.6755}\pm 0.14$ $\bm{{\color[rgb]{0,0,0}0.6386}\pm 0.14}$ ${\color[rgb]{0,0,0}0.5875}\pm 0.13$ $\bm{{\color[rgb]{0,0,0}0.5189}\pm 0.16}$ ${\color[rgb]{0,0,0}0.7329}\pm 0.15$ ${\color[rgb]{0,0,0}9.1902}\pm 6.40$ ${\color[rgb]{0,0,0}162.1719}\pm 211.33$ ${\color[rgb]{0,0,0}0.1054}\pm 0.04$
MOMENT(Goswami et al., 2024) ${\color[rgb]{0,0,0}0.5456}\pm 0.15$ ${\color[rgb]{0,0,0}0.6691}\pm 0.15$ ${\color[rgb]{0,0,0}0.6059}\pm 0.15$ ${\color[rgb]{0,0,0}0.5603}\pm 0.15$ ${\color[rgb]{0,0,0}0.4902}\pm 0.19$ ${\color[rgb]{0,0,0}0.7241}\pm 0.16$ $\bm{{\color[rgb]{0,0,0}8.5742}\pm 6.09}$ $\bm{{\color[rgb]{0,0,0}139.6036}\pm 168.57}$ $\bm{{\color[rgb]{0,0,0}0.0988}\pm 0.05}$
Supervised 1D ResNet-26 ${\color[rgb]{0,0,0}0.4623}\pm 0.16$ ${\color[rgb]{0,0,0}0.6212}\pm 0.17$ ${\color[rgb]{0,0,0}0.4684}\pm 0.18$ ${\color[rgb]{0,0,0}0.4944}\pm 0.14$ ${\color[rgb]{0,0,0}0.4025}\pm 0.15$ ${\color[rgb]{0,0,0}0.6224}\pm 0.18$ ${\color[rgb]{0,0,0}17.2733}\pm 18.49$ ${\color[rgb]{0,0,0}580.3867}\pm 961.79$ ${\color[rgb]{0,0,0}0.1803}\pm 0.15$
Rand Forest ${\color[rgb]{0,0,0}0.4152}\pm 0.10$ ${\color[rgb]{0,0,0}0.5911}\pm 0.15$ ${\color[rgb]{0,0,0}0.4671}\pm 0.12$ ${\color[rgb]{0,0,0}0.4337}\pm 0.10$ ${\color[rgb]{0,0,0}0.3328}\pm 0.08$ ${\color[rgb]{0,0,0}0.5977}\pm 0.09$ ${\color[rgb]{0,0,0}12.5655}\pm 7.70$ ${\color[rgb]{0,0,0}268.2220}\pm 283.46$ ${\color[rgb]{0,0,0}0.1495}\pm 0.08$
6.2.1. Background
To evaluate the effectiveness of Pulse-PPG, we compare it against additional baseline models. We include general-purpose time series foundation models, which are widely adopted for benchmarking due to their large-scale and diverse pretraining datasets, in this comparison. Specifically, we benchmark against Chronos (Ansari et al., 2024) and MOMENT (Goswami et al., 2024), two well-known general-purpose time series foundation models in this space. Chronos trains language models with suitable adaptations to time series data from domains that include energy, finance, economics, nature, retail, and transportation. MOMENT uses reconstruction of masked embeddings as a pre-training task, on datasets covering healthcare, human physiology, nature, power, economics, traffic, and weather. For supervised baselines, from classical machine learning, we use Random Forest, while from deep learning, we employ 1D ResNet-26, the same backbone used in the original Pulse-PPG model.
6.2.2. Experimental Design:
We train a Random Forest model with statistical features—mean, median, maximum, minimum, 25th percentile, and 75th percentile, computed from raw PPG windows corresponding to ground truth from respective downstream datasets. To train the ResNet model, we append a dense layer with output units equal to the number of classes, followed by an activation layer. For regression tasks, we use a dense layer with a single output unit without activation. Chronos (200 million) and MOMENT (385 million) produce embeddings of size 768 and 1024, respectively (compared to Pulse-PPG’s 28.5 million parameters with an embedding size of 512). We extract embeddings from these models for respective input windows from each of the downstream datasets and apply the same linear probe pipeline for classification and regression tasks.
6.2.3. Results:
We present downstream task performance of Pulse-PPG against supervised baselines and general-purpose time series models in Table 6 in the Appendix. We summarize the results with their average task-relevant (classification/regression) performance metrics in Table 2. Pulse-PPG was trained on a single dataset in a self-supervised fashion, yet it demonstrates balanced generalizability across diverse downstream tasks with a simple linear probe evaluation. Whereas, 1D ResNet-26 and Random Forest were trained separately on each downstream dataset. While ResNet-26 achieves exceptional performance for certain tasks, e.g., an F1-score of 0.7903 for PPG-DaLiA multiclass activity detection, its performance in multiclass stress classification for WESAD with an F1-score of 0.1998 is much worse. Pulse-PPG outperforms the Random Forest model on the majority of tasks.
General-purpose time series models benefit from exposure to diverse datasets, both in terms of nature and scale. This diversity enables them to learn periodicities and trends across contexts, helping them capture complex temporal patterns. Hence, although general-purpose time series models may outperform modality-specific foundation models like Pulse-PPG and PaPaGei in certain tasks, e.g., multi-class activity detection, the value of pretraining on noisy, real-world field PPG data is evident. As shown in Table 2, Pulse-PPG consistently ranks among the top two across a wide range of tasks focused on diverse applications in the physiological domain.
6.3. Experiment 3: Assessing Field-to-Lab Generalizability
6.3.1. Background:
Traditionally, mHealth models were built and tested in controlled lab settings due to a lack of diverse field datasets. As more real-world data became available, researchers began applying these lab-trained models to field data. However, we hypothesize that pre-training foundation models directly on diverse field data will yield better generalizability and performance.
6.3.2. Experimental Design
In these experiments, we assess how pre-training on wearable field PPG compares to pre-training on clean PPG. Similar to the prior experiments in Section 6.1, we utilize a linear probe for evaluation to ensure fair comparisons. From the prior experiments, we already have our Pulse-PPG pre-trained on field PPGs and PaPaGei pre-trained on the clinical PPGs.
In order to construct a large clean clinical PPG pre-training dataset for re-training Pulse-PPG, we utilize the curated MIMIC-III PPG waveform dataset from (Xu et al., 2022). This dataset is composed of 151,738 100 Hz 5-minute-long clean clinical PPGs from 18,210 patients. [^4] Unfortunately, PaPaGei did not release their full curation code for constructing their clinical pre-training dataset, making it impossible for us to train on the same clinical data. However, the MIMIC-III waveform did form a significant portion of their dataset. We then use this clean clinical PPG dataset to pre-train our Pulse-PPG model, following the same training procedures and hyperparameters that we used when we were pre-training on the MOODS field PPG dataset.
Next, we pre-train PaPaGei from scratch with the MOODS field dataset, using its original training procedure from their codebase located here: https://github.com/Nokia-Bell-Labs/PaPaGei-foundation-model.
[…]
This figure compares the generalizability of models pre-trained on wearable field data versus clinical data. It shows classification performance via F1 score and regression accuracy via MAPE, highlighting better field-to-lab generalizability from wearable data.
Figure 4. Comparison of Pre-training on Wearable Field data vs. Clinical Data to Assess Field-to-Lab Generalizability. The two plots show the relative performance of a linear probe evaluation for each model, with the left for classification via F1 score and the right for regression via Mean Average Percentage Error. 10/11 of the task data points reside above the slope, demonstrating how pre-training on wearable field data is highly generalizable, demonstrating strong field-to-lab generalizability and even field-to-clinical generalizability. This is contrary to the current standard line of research, which tends to focus on lab-to-field methods. See further discussion in Section 6.3.3.
6.3.3. Results
Figure 4 visualizes the difference in performance between Pulse-PPG pre-trained on the wearable field vs. clinical PPGs. Full results, including comparisons of PaPaGei pre-training with wearable field vs. clinical data, can be found in Table 7 in the Appendix, while referencing the second row named ”Pre-Train Data”.
Pulse-PPG consistently exhibits stronger performance on 10 out of 11 tasks when pre-trained on the Field Wearable PPG data versus when pre-trained on the clean Clinical PPG. For all other tasks except for sleep disturbance, the field wearable pre-training yielded consistently stronger results, sometimes substantially so. For example, for Field Wearable datasets, the F1 score for stress improved by .0848, and the F1 score for activity improved by .1218. Even for the clinical tasks, we see that pre-training on the wearable PPGs achieves much stronger performance (i.e., .10 F1 improvement in hypertension prediction and $\sim 1$ MAE improvement in systolic and diastolic BP regression). In some cases, such as in the WESAD evaluation datasets, one metric was worse, but all of the remaining metrics were stronger.
Pre-training Pulse-PPG on field wearable data appears to yield a more robust and generalizable representation, outperforming clinical pre-training, even on clinical tasks. Although performance gains in clinical tasks are modest, the consistency demonstrates that pre-training on field wearable data performs as well as, and sometimes better than, pre-training on clinical data. As discussed earlier in Section 6.1.3, relatively poor results on sleep can be explained by the lack of sleep data present in the MOODS dataset.
This trend of stronger performance when trained on wearable data is also observed in PaPaGei, though to a lesser extent. When pre-trained on wearable data, PaPaGei outperforms clinical pre-training in 6 out of 11 tasks, including five wearable PPG evaluation datasets (Activity - MOODS, Instant HR - PPG-DaLiA, Activity - PPG-DaLiA, Stress (2) - WESAD, and Stress (4) - WESAD). This aligns with the idea that pre-training and evaluating on wearable data introduces less domain shift. The difference in performance across pre-training settings between Pulse-PPG and PaPaGei can be attributed to their pre-training strategies. PaPaGei relies on PPG-specific morphologies, which may not generalize well across domains with varying PPG patterns (e.g., wearable vs. clinical). In contrast, Pulse-PPG leverages a learnable motif-similarity distance, enabling it to capture more general PPG patterns, which are applicable across different settings.
6.4. Experiment 4: Evaluation of Pulse-PPG Use-cases
6.4.1. Background
In this section, we quantitatively measure the effectiveness of Pulse-PPG for 11 downstream tasks, to provide further insight into its utility to the research community.
6.4.2. Experimental Design
We evaluate downstream performance using two standard evaluation methodologies:
  • Linear Probe: The model’s embeddings are frozen, and a logistic classifier or linear regression is trained on the embedding vectors, consistent with prior foundation model work (Abbaspourazad et al., 2023; Devlin et al., 2018; Oquab et al., 2023). This is a simple evaluation designed to evaluate the quality of the learned representations to generalize to new tasks. By requiring minimal computational resources, it also enables a broader range of researchers to utilize the model’s learned representations. Introducing a more sophisticated traditional-ML-based probe would confound the results, with it being more difficult to isolate representation quality from the strength of the classifier.
  • Fine Tune: This evaluation is designed to boost task-specific performance by training the entire network end-to-end to adapt to a specific task. This highlights the adaptability of the foundation model, showing its potential to serve as a strong starting point for specialized tasks. While computationally intensive, it is essential for applications requiring peak performance.
Table 3. Pulse-PPG performance across wearable and clinical PPG tasks. There are three evals: Naive, predicting majority class or mean; Linear Probing, which trains a simple head on the Pulse-PPG’s frozen embeddings; and Fine-Tuning, which trains Pulse-PPG’s network end-to-end.
Pulse-PPG
Eval Method: Naive LP FT
Wearable Lab Instant HR (R) PPG-DaLiA MAE 19.34 8.936 3.705
MSE 522.7 149.3 59.81
MAPE 0.2742 0.1166 0.0484
Activity (9) PPG-DaLiA F1 0.0455 0.3039 0.3648
Accuracy 0.2572 0.4104 0.4234
Precision 0.0286 0.3882 0.4398
Recall 0.1111 0.3107 0.3663
AUPRC 0.0728 0.3504 0.4056
AUROC 0.5000 0.8051 0.8246
Stress (2) WESAD F1 0.4065 0.8759 0.9392
Accuracy 0.6849 0.8904 0.9452
Precision 0.3425 0.8688 0.9259
Recall 0.5000 0.8848 0.9600
AUPRC 0.3151 0.9410 0.9790
AUROC 0.5000 0.9687 0.9852
Stress (4) WESAD F1 0.1496 0.5431 0.5582
Accuracy 0.4270 0.6517 0.7079
Precision 0.1067 0.6003 0.5515
Recall 0.2500 0.5716 0.5812
AUPRC 0.2584 0.5671 0.5898
AUROC 0.5000 0.7807 0.7630
Wearable Field Stress (2) MOODS F1 0.3859 0.5398 0.5872
Accuracy 0.6285 0.6156 0.6247
Precision 0.3143 0.5606 0.5911
Recall 0.5000 0.5460 0.5859
AUPRC 0.3715 0.4275 0.5729
AUROC 0.5000 0.5691 0.5909
Activity (2) MOODS F1 0.4574 0.6859 0.7992
Accuracy 0.8430 0.8705 0.8930
Precision 0.4215 0.7835 0.7974
Recall 0.5000 0.6509 0.8010
AUPRC 0.1570 0.5835 0.8596
AUROC 0.5000 0.8722 0.9196
Clinical Sleep Disturbance (2) SDB F1 0.3945 0.4630 0.5782
Accuracy 0.6515 0.5364 0.6428
Precision 0.3258 0.4633 0.5911
Recall 0.5000 0.4673 0.5772
AUPRC 0.3485 0.3085 0.5904
AUROC 0.5000 0.4437 0.6243
Systolic BP (R) PPG-BP MAE 17.21 13.62 12.33
MSE 455.9 286.5 280.7
MAPE 0.1345 0.1065 0.0939
Diastolic BP (R) PPG-BP MAE 10.12 8.878 8.695
MSE 147.9 118.2 119.8
MAPE 0.1407 0.1232 0.1209
Average HR (R) PPG-BP MAE 7.965